On Martin's car the total length of the windshield wiper arm, plus the blade, is 31 inches and it turns through a 110 degree angle. If the blade alone is 18 inches long to the nearest square inch how much area of the windshield is cleaned by the wiper blade?
Answer by joey_quizan(15)

Sol:
A=Total Area
A=2*31*3.1461*(110/180)=119 sq.in.
a=area cleaned by wiper
a= 119-2*(31-18)*3.1461*(110/180)
Ans: a= 69 sq.in.